Two complementary angles are ​(77x-11​)degrees° and ​(44x+33​)degrees°Find their measures.
​(Simplify your answer. Type an integer or a​ fraction.)

Answer by Boreal(15235)

They add to 90 degrees
77x-11+44x+33=90
121x+22=90
121x=68
x=68/121
If need a fraction, then the first angle is (77*68/121)-11=32 3/11 degrees
The second angle would be (44*68/121)+33=57 8/11 degrees. They add to 89 11/11 degrees or 90 degrees.